OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)? A) y = 4x − 2 B) y = 2x − 4 C) y = −3(x − 4) D) y = 3(x + 4)

OpenStudy (anonymous):

1- find the slope using the 2 points

OpenStudy (anonymous):

\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\left( x-x _{1} \right)\]

OpenStudy (anonymous):

2- use any point to find the equation of the line y-y1=m(x-x1) m: is the slope x1 and y1 is the point